OpenStudy (anonymous):

Find the partial derivatives of the function. w = sqrt(r^2+s^2+t^2) There are 3 answers: dw/dr=? dw/ds=? dw/dt=?

OpenStudy (kainui):

Show me your best guess and we'll go from there.

OpenStudy (anonymous):

I don;t know anything about partial derivatives, can you please show me how to do it?

OpenStudy (kainui):

Well if you know how to take normal derivatives, you should be perfectly fine here. The only difference is that you've probably just locked yourself into thinking derivatives are only with respect to x. So just pretend when it says dw/dr that when you see "r" it's like an "x" and everything else you see is a constant. That means s and t. A quick example: d/dr of w=(s^2)(r^3)/t is dw/dr=3(r^2)*(s^2)/t See how everything else was constant and just turned the r^3 into 3r^2? It's really that easy.

OpenStudy (anonymous):

So do I just take the derivative in terms of r for the first one?

OpenStudy (kainui):

Yep

OpenStudy (kainui):

give it a shot.

OpenStudy (anonymous):

Oh I got it thanks.

OpenStudy (kainui):

Cool, want to show me an example of one you did to make sure, or are you good?

OpenStudy (anonymous):

No thanks I know I got it right because there's instant feedback.